# The Unreasonable Effectiveness of Model Merging for Cross-Lingual Transfer in LLMs

### v1

@@ Abstract
Large language models (LLMs) still struggle across tasks outside of high-resource languages. In this work, we investigate cross-lingual transfer to lower-resource languages where task-specific post-training data is scarce. Building on prior work, we first validate that the subsets of model parameters that matter most for mathematical reasoning and multilingual capabilities are distinctly non-overlapping. To exploit this implicit separability between task and target language parameterization, we develop and analyze numerous modular frameworks to improve the composition of the two during fine-tuning. These methods generally employ freezing parameters or post hoc model merging to assign math and language improvement to different key parts of the LLM. In the absence of in-language math data, we demonstrate that the modular approaches successfully improve upon baselines across three languages, four models, and two fine-tuning paradigms (full and LoRA). Furthermore, we identify the most consistently successful modular method to be fine-tuning separate language and math experts and model merging via Layer-Swapping (Bandarkar et al., 2025), somewhat surprisingly. We offer possible explanations for this result via recent works on the linearity of task vectors. We further explain this by empirically showing that reverting less useful fine-tuning updates after training often outperforms freezing them from the start.
@@ 1 Introduction
Figure 1: Illustration of the three methods that induce modularity by imposing target language capabilities (brown) and mathematical reasoning (blue) on separate LLM parameters. [1] is from Bandarkar et al. (2025)
@@
With these datasets, we evaluate numerous training paradigms that incentivize the model, to varying degrees, to learn multilingual or math capabilities in specific parameters. We organize the settings along two axes: (1) whether the models are optimized separately or together over the two SFT datasets and (2) whether the same or separate model parameters are trained on the datasets. When the models are trained separately, we combine the learned capabilities using model merging methods such as variants of Layer-Swapping (Bandarkar et al., 2025). To train separate model parameters, we start by dividing all parameters into two partitions according to prior work: (1) one set allocated to target language training and (2) one set to English math. Only allocated parameters are fine-tuned, while the opposite partition is frozen. We additionally develop a method to train separate parameters in a single, joint training by frequently freezing and unfreezing parameters to simulate simultaneous training.
@@
Amongst our modular solutions, we surprisingly find that post hoc model merging via Layer-Swapping outperforms more coordinated multi-task fine-tuning approaches. To contextualize this counterintuitive result, we explore recent academic literature that help explain the phenomenon. We provide empirical evidence for training all model parameters, even if large portions will be discarded during Layer-Swapping. While these subsets of task vectors are unproductive, freezing them during fine-tuning leads to less optimal updates to the target parameters. Notably, we rationalize that the fine-tuning task vectors ($\Delta$s) are quite linear within individual parameter blocks (Dai et al., 2025), meaning they can be added, scaled, or interpolated as linear components (Adilova et al., 2024).
@@ 2.1 Cross-Lingual Transfer
However, cross-lingual alignment in more modern decoder-only models has become less methodical because of the lack of universal embedding (Kargaran et al., 2024). Since most popular LLMs have been trained on a majority English corpora, recent works have examined how much intrinsic cross-lingual transfer occurs at different training stages (Choenni et al., 2023; Wang et al., 2024). These large models have broader generalization and robustness, but still fail to transfer much of their capabilities across languages (Philippy et al., 2023). Recent works have identified prompting methods (Shi et al., 2023; Zhang et al., 2024) or post-training data augmentation (Dang et al., 2024a; She et al., 2024; Lai et al., 2024) to help generalization.
@@ 2.2 Modularity in Multilingual NLP
Even in dense LLMs, however, interpretability research has identified the emergence of effective modularity (Csordás et al., 2021) as LLM parameters scale (Zhang et al., 2022; Qiu et al., 2024). Principally, numerous recent works have identified that just a few transformer layers at the top and bottom of English-centric LLMs are responsible for multilingual capabilities, notably by mapping input and output into a universal representation (Kojima et al., 2024; Wendler et al., 2024; Tang et al., 2024b; Alabi et al., 2024).
@@ 2.3 Model Merging
Model merging is the practice of combining the weights of multiple checkpoints of the same model architecture into a singular model. While averaging models is a fundamental machine learning approach to increase statistical robustness (Breiman, 1996), the averaging of model checkpoints, dubbed a model soup by Wortsman et al. (2022), has re-emerged in large-scale LLMs as a method to increase model robustness. More importantly, it also increases the search space of valid model variants at any given training step without additional costly training runs (Llama et al., 2024). However, simple weight averaging is vulnerable to negative transfer, or interference, between checkpoints so numerous methods have been presented to selectively merge parameters (Ilharco et al., 2023a; Yadav et al., 2023; Yu et al., 2024a). Surprisingly, training models on separate data and then merging can often outperform a single training run on mixed data (Tang et al., 2024a; Aakanksha et al., 2024) and has shown to be highly effective in large-scale multilingual pretraining (Dang et al., 2024b). For cross-lingual transfer in particular, Ansell et al. (2022) showed that sparse fine-tuning can lead to better composition. Bandarkar et al. (2025) extended this by notably identifying that mathematical reasoning was concentrated in parameters different from multilingual capabilities. As a result, model variants trained on English math data and multilingual data can be combined by Layer-Swapping, or swapping the transformer layers most important to each.
@@ 3.3 Parameter Allocation
To determine which parameters to “allocate” to each capability, we rely on a mix of interpretability papers and small-scale empirical tests. As mentioned in Section 2.2, numerous papers have identified the most important parameters for multilingual capabilities to be the first few and last few transformer layers of LLMs. These works, however, typically discuss mostly English-centric models (such as Llama 3.1 and Falcon 3). We therefore need to evaluate this for bilingual and multilingual models like Qwen2.5 and Aya Expanse. For mathematical reasoning, we note that Bandarkar et al. (2025) identifies the middle and late-middle transformer layers as being the most important. This work, and numerous others (Voita et al., 2019; Ma et al., 2021; Zhao et al., 2024), similarly identifies multi-head attention parameters as critical to multilingual capabilities, as opposed to multi-layer perceptron parameters.
@@ 3.4 Training
For SFT data, we create four datasets, one for math in English and one instruction dataset for each of the three target languages. The math instruction dataset consists of English math word problems from the Orca-Math synthetic dataset (Mitra et al., 2024). For the language datasets, we replicate the creation of “generic” instruction fine-tuning datasets from Bandarkar et al. (2025) by combining samples from open-source instruction and task-specific datasets. Importantly, there are no math samples in these multi-task language datasets. We provide specific details and citations for these data collections in Appendix A.6.
@@ 4.3 Layer-Swapping (Separate Training, Same Parameters)
For this setting, we exactly recreate the method presented by Bandarkar et al. (2025). Starting from the same base model, separate variants are trained on different tasks, dubbed “experts”. Concretely, one expert has been trained on the English math data, and the other on the target language instruction dataset. To recompose a single model, the top and bottom transformer layers from the target language expert replace those in the math expert, while the math experts’ middle layers remain. We additionally implement the equivalent of this methodology with LoRA, where the set of adapters is merged by combining the adapters corresponding to parameters that would be swapped. Note that we do not retrain these experts and simply use the checkpoints from our baseline trainings.
@@ 6 Discussion
We seek to explain why concatenating transformer layers from separately fine-tuned “experts" is so seamless. Task vectors (Ilharco et al., 2023b) are the $\Delta$s that result from fine-tuning (i.e., $\theta_{FT}-\theta_{0}$). Task vector linearity refers to the property that linear combinations of such task vectors form a coherent, effective model. Ortiz-Jimenez et al. (2023) identifies that linearized task vectors exhibit better mergeability. Meanwhile, when fine-tuning heavily post-trained models like those used in our experiments, recent works show that updates to individual model layers exhibit significant linearity (Zhou et al., 2024; Razzhigaev et al., 2024; Dai et al., 2025). Furthermore, research on mode connectivity (Frankle et al., 2020; Garipov et al., 2018) shows individual transformer layers can be smoothly interpolated (Zhou et al., 2023; Adilova et al., 2024). These works provide explanation for why ad hoc Layer-Swapping is not more degradative.
@@ A.6 SFT Datasets
xP3 MT dataset from BigScience, with FLoRes samples removed (Muennighoff et al., 2022) https://huggingface.co/datasets/bigscience/xP3mt

### v2

@@ Abstract
Large language models (LLMs) still struggle across tasks outside of high-resource languages. In this work, we investigate cross-lingual transfer to lower-resource languages where task-specific post-training data is scarce. Building on prior work, we first validate that the subsets of model parameters that matter most for mathematical reasoning and multilingual capabilities are distinctly non-overlapping. To exploit this implicit separability between task and target language parameterization, we develop and analyze numerous modular frameworks to improve the composition of the two during fine-tuning. These methods generally employ freezing parameters or post hoc model merging to assign math and language improvement to different key parts of the LLM. In the absence of in-language math data, we demonstrate that the modular approaches successfully improve upon baselines across three languages, four models, and two fine-tuning paradigms (full and LoRA). Furthermore, we identify the most consistently successful modular method to be fine-tuning separate language and math experts and model merging via Layer-Swapping (Bandarkar et al., 2025a), somewhat surprisingly. We offer possible explanations for this result via recent works on the linearity of task vectors. We further explain this by empirically showing that reverting less useful fine-tuning updates after training often outperforms freezing them from the start.
@@ 1 Introduction
Figure 1: Illustration of the three methods that induce modularity by imposing target language capabilities (brown) and mathematical reasoning (blue) on separate LLM parameters. [1] is from Bandarkar et al. (2025a)
@@
With these datasets, we evaluate numerous training paradigms that incentivize the model, to varying degrees, to learn multilingual or math capabilities in specific parameters. We organize the settings along two axes: (1) whether the models are optimized separately or together over the two SFT datasets and (2) whether the same or separate model parameters are trained on the datasets. When the models are trained separately, we combine the learned capabilities using model merging methods such as variants of Layer-Swapping (Bandarkar et al., 2025a). To train separate model parameters, we start by dividing all parameters into two partitions according to prior work: (1) one set allocated to target language training and (2) one set to English math. Only allocated parameters are fine-tuned, while the opposite partition is frozen. We additionally develop a method to train separate parameters in a single, joint training by frequently freezing and unfreezing parameters to simulate simultaneous training.
@@
[…]
@@ 2.1 Cross-Lingual Transfer
However, cross-lingual alignment in more modern decoder-only models has become less methodical because of the lack of universal embedding (Kargaran et al., 2025). Since most popular LLMs have been trained on a majority English corpora, recent works have examined how much intrinsic cross-lingual transfer occurs at different training stages (Choenni et al., 2023; Wang et al., 2024). These large models have broader generalization and robustness, but still fail to transfer much of their capabilities across languages (Philippy et al., 2023). Recent works have identified prompting methods (Shi et al., 2023; Zhang et al., 2024) or post-training data augmentation (Dang et al., 2024a; She et al., 2024; Lai et al., 2024) to help generalization.
@@ 2.2 Modularity in Multilingual NLP
Even in dense LLMs, however, interpretability research has identified the emergence of effective modularity (Csordás et al., 2021) as LLM parameters scale (Zhang et al., 2022; Qiu et al., 2024; Chen et al., 2025). Principally, numerous recent works have identified that just a few transformer layers at the top and bottom of English-centric LLMs are responsible for multilingual capabilities, notably by mapping input and output into a universal representation (Kojima et al., 2024; Wendler et al., 2024; Tang et al., 2024b; Alabi et al., 2024; Wu et al., 2025). Similar patterns are observed in modern sparse mixture-of-experts LLMs, where it is also observed that language-specialized experts are completely distinct from task/domain-specialized ones (Bandarkar et al., 2025b).
@@ 2.3 Model Merging
Model merging is the practice of combining the weights of multiple checkpoints of the same model architecture into a singular model. While averaging models is a fundamental machine learning approach to increase statistical robustness (Breiman, 1996), the averaging of model checkpoints, dubbed a model soup by Wortsman et al. (2022), has re-emerged in large-scale LLMs as a method to increase model robustness. More importantly, it also increases the search space of valid model variants at any given training step without additional costly training runs (Llama et al., 2024). However, simple weight averaging is vulnerable to negative transfer, or interference, between checkpoints so numerous methods have been presented to selectively merge parameters (Ilharco et al., 2023a; Yadav et al., 2023; Yu et al., 2024a). Surprisingly, training models on separate data and then merging can often outperform a single training run on mixed data (Tang et al., 2024a; Aakanksha et al., 2024) and has shown to be highly effective in large-scale multilingual pretraining (Dang et al., 2024b). For cross-lingual transfer in particular, Ansell et al. (2022) showed that sparse fine-tuning can lead to better composition. Bandarkar et al. (2025a) extended this by notably identifying that mathematical reasoning was concentrated in parameters different from multilingual capabilities. As a result, model variants trained on English math data and multilingual data can be combined by Layer-Swapping, or swapping the transformer layers most important to each.
@@ 3.3 Parameter Allocation
To determine which parameters to “allocate” to each capability, we rely on a mix of interpretability papers and small-scale empirical tests. As mentioned in Section 2.2, numerous papers have identified the most important parameters for multilingual capabilities to be the first few and last few transformer layers of LLMs. These works, however, typically discuss mostly English-centric models (such as Llama 3.1 and Falcon 3). We therefore need to evaluate this for bilingual and multilingual models like Qwen2.5 and Aya Expanse. For mathematical reasoning, we note that Bandarkar et al. (2025a) identifies the middle and late-middle transformer layers as being the most important. This work, and numerous others (Voita et al., 2019; Ma et al., 2021; Zhao et al., 2024), similarly identifies multi-head attention parameters as critical to multilingual capabilities, as opposed to multi-layer perceptron parameters.
@@ 3.4 Training
For SFT data, we create four datasets, one for math in English and one instruction dataset for each of the three target languages. The math instruction dataset consists of English math word problems from the Orca-Math synthetic dataset (Mitra et al., 2024). For the language datasets, we replicate the creation of “generic” instruction fine-tuning datasets from Bandarkar et al. (2025a) by combining samples from open-source instruction and task-specific datasets. Importantly, there are no math samples in these multi-task language datasets. We provide specific details and citations for these data collections in Appendix A.6.
@@ 4.3 Layer-Swapping (Separate Training, Same Parameters)
For this setting, we exactly recreate the method presented by Bandarkar et al. (2025a). Starting from the same base model, separate variants are trained on different tasks, dubbed “experts”. Concretely, one expert has been trained on the English math data, and the other on the target language instruction dataset. To recompose a single model, the top and bottom transformer layers from the target language expert replace those in the math expert, while the math experts’ middle layers remain. We additionally implement the equivalent of this methodology with LoRA, where the set of adapters is merged by combining the adapters corresponding to parameters that would be swapped. Note that we do not retrain these experts and simply use the checkpoints from our baseline trainings.
@@ 6 Discussion
[…]
@@ A.6 SFT Datasets
xP3 MT dataset from BigScience, with FLoRes samples removed (Muennighoff et al., 2023) https://huggingface.co/datasets/bigscience/xP3mt